The density of a mixture of O2 and N2 at STP is 1.3g/L. Calculate the partial pressure of O2

Grade:11

1 Answers

Arun
6 years ago
 

given

 

N2 and O2 mixture    i.e N2+O2      density = 1.3g/l at stp

 

 

 

i.e     1 lit = 1.3g

 

therefore   22.4 lit = 22.4 * 1.3

 

     = 29.12 g

 

we know that at 22.4 lit the weight is known as molecular weight

 

therfore mw(M)= 29.12 g

 

wkt.,n=w/M

 

      w= n*M

 

for given eq., N2 +O2

 

   total weight at 22.4 lit (stp) is = 28x + 32 (1-x)

 

=>   29.12  =   28x+32-32x

 

therefore   x = 0.72

 

therefore  no.of moles of N2= 0.72 and O2 = 1- 0.72 = 0.28

 

therefore partial pressure of O2 at stpPO2 = n O2/(n O2 + n N2 )  *  p total

 

                                                         p O2    =   0.28/(0.28+0.72)   *    1atm

 

                                                                     p  O2 =    0.28 atm.